# A Comprehensive Survey on Graph Reduction: Sparsification, Coarsening, and Condensation

### v1

@@ 3.2.2 Reconstruction-Free Methods
where $\mathcal{V}^{\prime}$ denotes the selected index from $\mathcal{V}$, $\bar{\mathcal{V}^{\prime}}=\mathcal{V}-\mathcal{V}^{\prime}$, and ${\bf L}_{\mathcal{V},\mathcal{V}^{\prime}}$ is the submatrix of ${\bf L}$ whose row index is $\mathcal{V}$ and column index is $\mathcal{V}^{\prime}$. Recently,  Sugiyama and Sato (2023) extend it to a directed graph with self-loop. Some methods attempt to find the most informative summaries of a network. To analyze social networks with diverse attributes and relations, SNAP (Tian et al., 2008) produces a summary graph where every node inside a super node has the same values for selected attributes, and is adjacent to nodes with the same selected relations. k-snap (Tian et al., 2008) relaxes this homogeneity requirement and allows users to control the resolutions of summaries. AGSUMMARY (Wu et al., 2014) utilizes the Minimum Description Length principle to design a cost function and compute an optimal summary by neighborhood greedy strategy. Since the former two methods only apply to discrete attributes, CANCEL (Zhang et al., 2010) extends the node attributes to continual ones with adaptive cutoffs and proposes a comprehensive metric named interestingness. To extend the above methods focusing on only one task, CoarseNet (Purohit et al., 2014) tries to find a succinct representation of the original network that preserves important diffusive characters, which can be applied to influence maximization and detecting patterns of propagation. Netgist (Amiri et al., 2018) defines a task-based graph summarization problem and uses RL to create a flexible framework for learnable node merging policies.
3.2.3 Remarks on Graph Coarsening
Graph Coarsening in GNNs. There are growing numbers of works that combine coarsening with GNNs. For instance, SCAL (Huang et al., 2021) first trains a GNN model in a graph coarsened by LV, with super node label defined as ${\bf Y}^{\prime}=\operatorname{argmax}({\bf P}{\bf Y})$ and then directly uses this model to inference. Following the same framework, Generale et al. (2022) define the relation summarization and uses R-GCN (Schlichtkrull et al., 2018) for knowledge graphs. To mimic the pooling layer in the convolutional neural network, Such et al. (2017) make the mapping matrix learnable and produce a pooled graph reduced to fewer nodes layer by layer. CONVMATCH (Dickens et al., 2023) merges nodes that are equivalent or similar w.r.t. the GCN convolution operation. Similarly,  Buffelli et al. (2022) match the node embeddings output by GNNs among graphs in different coarsening ratios. VNG (Si et al., 2023) highlights the ongoing challenge of efficiently deploying a GNN model in online applications when connections exist between testing nodes and training nodes. To tackle this issue, they match the forward propagation by applying weighted k-means to obtain the mapping matrix ${\bf C}$.
Connections with Graph Clustering/Partition. Partition and clustering in graphs are long-developed areas. Graph partition aims to find a split for a graph with the least cost, e.g., cutting the fewest edges. The representative partition method, METIS (Karypis and Kumar, 1997), coarsens a graph iteratively by edge contraction, splits the nodes in the coarsest graph, and reversely maps them to the original graph. This reconstruction-based framework is widely used for partition (Safro et al., 2015), which means the areas of partition and coarsening are mutually reinforcing. Graph clustering attempts to find groups of nodes with high in-group edge density, and relatively low out-group density (Tsitsulin et al., 2023). Thus, by mapping these dense groups to super nodes and aggregating them, any clustering method can apply to the coarsening strategy.
@@ 3.2.3 Remarks on Graph Coarsening
DisCo (Xiao et al., 2024) Disentangled Condensation NC X, A, Y
@@ 3.3.1 Matching-Based Methods
Several subsequent studies target at improving GM for graph condensation to enhance the effectiveness of GCond (Gao and Wu, 2023; Yang et al., 2023; Feng et al., 2023; Mao et al., 2023; Li et al., 2023b; Gao et al., 2023; Anonymous, 2024). Unlike GCond, which uses a single fully connected graph to generate the condensed graph dataset $\mathcal{S}$, MSGC (Gao and Wu, 2023) is introduced to leverage multiple sparse graphs to create diverse neighborhoods for nodes that enhance the capturing of neighborhood information. This, in turn, allows GNNs to generate more informative embeddings in the condensed graphs. Regarding the generalizability of GCond across different GNN architectures, SGDD (Yang et al., 2023) is proposed to explicitly prevent overlooking the original graph dataset structure $\mathbf{A}$ by broadcasting it into the construction of synthetic graph structure $\mathbf{A^{\prime}}$. In this way, it is shown that SGDD reduces the Laplacian Energy Distribution (LED) (Das et al., 2016; Gutman and Zhou, 2006) shift crossing various datasets significantly compared to GCond. In addition to the node classification task, to validate the effectiveness of SGDD, extensive link prediction problems have been explored. Gao et al. (2023) identify the potential issues in existing graph condensation methods for inductive node representation learning and emphasize the under-explored need for an explicit mapping between original and synthetic nodes. Consequently, a GM-based method named MCond was introduced, which explicitly learns a sparse mapping matrix to smoothly integrate new nodes into the synthetic graph for inductive representation learning. MCond employs an alternating optimization scheme compared to GCond, allowing the synthetic graph and mapping matrix to take turns updating toward dedicated objectives. Furthermore, CTRL (Anonymous, 2024) argues the limited approach of using cosine similarity for gradient matching, leading to biases, and suggests adding gradient magnitudes into the objective function introduced in GCond for a more accurate match. Their empirical findings also show that this approach better aligns frequency distributions between condensed and original graphs.
@@
Despite the effectiveness of the previously mentioned graph condensation methods, Feng et al. (2023) recognize that these methods tend to exhibit fairness issues. By identifying the group fairness [^3] Group fairness in algorithms ensures unbiased and fair treatment across diverse demographic groups. It seeks to prevent any form of discrimination or bias against specific groups within the algorithmic decision-making process (Mehrabi et al., 2021), it demonstrated that as distillation performance increases, fairness (Demographic Parity $\Delta_{DP}$) decreases (Feng et al., 2023). Particularly, it is showcased that, by measuring the fairness of GNNs trained on original graphs versus those trained on condensed graphs, an improvement in performance correlates with heightened fairness issues in the synthetic condensed graph. To address this challenge, FGD was introduced, as a fair graph condensation method. This was achieved by incorporating the coherence metric into the GM loss function outlined in Eq. (8a). Particularly, the coherence metric is a bias calculator that captures the variance of the estimated sensitive group membership. Similarly, to address the fairness issue of current graph condensation methods, Mao et al. (2023) propose graph condensation with Adversarial Regularization (GCARe), which is a method that directly regularizes the condensation process to distill the knowledge of different subgroups fairly into resulting graphs.
@@
where ${\mathcal{C}}$ is the set of node classes, ${\boldsymbol{E}}_{c}$ and $\tilde{\boldsymbol{E}}_{c}$ are the embeddings of nodes with class $c$ in the original and condensed graph, respectively, and $r_{c}$ is the class ratio for class $c$. Likewise, GCDM (Liu et al., 2022) regards the original graph as a distribution of receptive fields and seeks to synthesize a smaller graph whose receptive fields share a similar distribution to that of the original graph. Other works, such as PUMA (Liu et al., 2023d) and FedGKD (Pan et al., 2023), employ a similar approach for various applications in continual learning and federated learning, respectively.
@@ 3.3.2 Kernel Ridge Regression Methods
where $\epsilon>0$ is a KRR hyper-parameter, $\mathbf{K}_{\mathcal{TS}}$ is the kernel matrix between original and synthetic graphs while $\mathbf{K}_{\mathcal{SS}}$ is the kernel matrix between synthetic graphs [^4] Each kernel indicates infinitely wide multi-layer GNNs trained by gradient descent through squared loss (Du et al., 2019). Also, $\mathbf{y}_{\mathcal{S}}$ and $\mathbf{y}_{\mathcal{T}}$ are the concatenated graph labels from real dataset and synthetic dataset, respectively. Also, Zheng et al. (2023) propose SFGC, a structure-free graph condensation method using KRR that only outputs the condensed node features $\mathbf{X}^{\prime}$, as the structure information of the real graphs is embedded in ${\bf X}^{\prime}$.
@@ 4.2 Continual Graph Learning
Continual learning  (De Lange et al., 2021) aims to learn on a stream of data from a stationary data distribution, which requires tackling the issue of catastrophic forgetting, i.e., new data can interfere with current knowledge and erase it. The common strategy is memory replay (Kemker and Kanan, 2018), which involves storing representative samples from previous tasks in a buffer to recall knowledge when needed. In the context of graphs, continual learning can be benefited by informative reduced graphs. As introduced in Section 3.3 CaT (Liu et al., 2023c) is applied to continual graph learning by condensing the incoming graph and updating the model with condensed graphs, not the whole incoming graph. To further improve CaT,  Liu et al. (2023d) introduce PUMA which utilizes pseudo-labeling to incorporate data from unlabeled nodes, boosting the informativeness of the memory bank and addressing the problem of neglected unlabeled nodes. In addition, the sparsification method (Zhang et al., 2023) reduces the number of nodes and edges according to the Ricci curvature evaluation and stores them into a replay buffer for continual learning.
@@ 4.3 Visualization & Explanation
The reduced dataset is not only easier to parse by algorithms and computers but also more friendly for people to understand, which leads to the application of visualization and explanation. For instance, Zhao et al. (2018) combine spectral graph coarsening method (Loukas and Vandergheynst, 2018) and sparsification methods (Feng, 2016) to develop a nearly linear time algorithm for multilevel graph visualization. Pyramid Transform (Shuman et al., 2015) selects nodes corresponding to top-$k$ eigenvector repeatedly and creates a multi-resolution view of large graphs. In addition, k-core decomposition has been widely used for graph visualization and finding specific structural characteristics of networks (Alvarez-Hamelin et al., 2005). As mentioned in Section 3.2, some coarsening methods like k-snap (Tian et al., 2008) and CANCEL (Zhang et al., 2010) are tailed for customized discovery, which means the user can set the granularity of the graph structure and then find perspectives they are interested in. Graph condensation is also used for visualization and explanation. For instance, GCond (Jin et al., 2022b) observes patterns from original data by visualizing the reduced graph; GDM (Nian et al., 2023) employs condensation to explain GNN behavior during the training process.
@@ 4.4 Privacy
It has been empirically investigated that reduced datasets offer good privacy preservation (Bonchi et al., 2014; Dong et al., 2022). However, the degree of reduction always influences the utility, i.e., preservation of key information. Dataset condensation can be a promising technique to solve the dilemma (Dong et al., 2022). Under a federated learning framework with particular concerns for local data privacy, FedGKD (Pan et al., 2023) leverages a features extractor that condenses a small graph in each round. Then, the local models are trained on the condensed graphs before federated aggregation.
@@ 4.5 Data Augmentation
Graph data augmentation (Zhao et al., 2022, 2021) is commonly used to enrich the data and improve the model performance. Methods for graph reduction can be employed to generate various perspectives of a graph by repeatedly applying reductions at different ratios, thereby augmenting the data for subsequent models. For example, HARP (Chen et al., 2018) coarsens a graph in a series of levels and then embeds the hierarchy of graphs from the coarsest one to the original. In each step, HARP initializes node embeddings in the finer-level graph using the mapped embeddings from super nodes in the coarser level. By employing various graphs in each iteration, this method augments the training data. Meanwhile, MILE (Liang et al., 2021) enhances this process by substituting the random walk employed in HARP with GNNs to improve the embedding quality and efficiency. DistMILE (He et al., 2021) advances the MILE framework by adopting high-performance computing techniques. As a condensation method, MSGC (Gao and Wu, 2023) initializes multiple small graphs by various connection schemes and employs gradient matching to optimize them. This process results in a variety of node embeddings sets, increasing diversity and thereby augmenting the data.
@@ 5.4 Distribution Shift
It is a common observation that GNNs often exhibit poor generalization on a test set when there exists a disparity between the distributions of the training and test sets (Wu et al., 2022a; Li et al., 2022). When GNNs are trained on reduced graphs and evaluated on graphs from the original distribution, a distribution shift may occur due to the reduction process, which eliminates substantial amounts of graph elements. However, consensus is lacking regarding the definition of graph data distribution or the selection of specific properties to represent the distribution accurately. While several condensation methods utilize one kind of distribution matching as we mentioned in Section 2, other measures of distribution may change after the reduction, e.g., size shift (Buffelli et al., 2022). Future graph reduction should consider the potential distribution shift issues and preserve properties related to distribution to enhance the generalization of models trained on reduced graphs.

### v4

@@ 3.2.3 Remarks on Graph Coarsening
where $\mathcal{V}^{\prime}$ denotes the selected index from $\mathcal{V}$, $\bar{\mathcal{V}^{\prime}}=\mathcal{V}-\mathcal{V}^{\prime}$, and ${\bf L}_{\mathcal{A},\mathcal{B}}$ is the submatrix of ${\bf L}$ whose row index is $\mathcal{A}$ and column index is $\mathcal{B}$. Recently,  Sugiyama and Sato (2023) extend it to a directed graph with self-loop. In contrast to selecting nodes arbitrarily,  Fang et al. (2010) calculate Schur complement after finding the largest node set consisting of nodes not adjacent to each other.
Connections with Graph Clustering/Partition. Partition and clustering in graphs are long-developed areas. Graph partition aims to find a split for a graph with the least cost, e.g., cutting the fewest edges. The representative partition method, METIS (Karypis and Kumar, 1997), coarsens a graph iteratively by pairwise contraction, splits the nodes in the coarsest graph, and reversely maps them to the original graph. This reconstruction-based framework is widely used for partition (Safro et al., 2015), which means the areas of partition and coarsening are mutually reinforcing. Graph clustering attempts to find groups of nodes with high in-group edge density, and relatively low out-group density (Tsitsulin et al., 2023). Thus, by mapping these dense groups to super nodes and aggregating them, any clustering method can apply to the coarsening strategy.
@@
SFGC (Zheng et al., 2023b) Trajectory Matching NC X, A, Y
GEOM (Zhang et al., 2024b) NC X, A, Y
@@ 3.3.1 Matching-Based Methods
Several subsequent studies target at improving GM for graph condensation to enhance the effectiveness of GCond (Gao and Wu, 2023; Yang et al., 2023; Feng et al., 2023; Mao et al., 2023; Li et al., 2023b; Gao et al., 2023; Zhang et al., 2024a). Unlike GCond, which uses a single fully connected graph to generate the condensed graph dataset $\mathcal{S}$, MSGC (Gao and Wu, 2023) is introduced to leverage multiple sparse graphs to create diverse neighborhoods for nodes that enhance the capturing of neighborhood information. This, in turn, allows GNNs to generate more informative embeddings in the condensed graphs. Regarding the generalizability of GCond across different GNN architectures, SGDD (Yang et al., 2023) is proposed to explicitly prevent overlooking the original graph dataset structure $\mathbf{A}$ by broadcasting it into the construction of synthetic graph structure $\mathbf{A^{\prime}}$. In this way, it is shown that SGDD reduces the Laplacian Energy Distribution (LED) (Das et al., 2016; Gutman and Zhou, 2006) shift crossing various datasets significantly compared to GCond. In addition to the node classification task, to validate the effectiveness of SGDD, extensive link prediction problems have been explored. Gao et al. (2023) identify the potential issues in existing graph condensation methods for inductive node representation learning and emphasize the under-explored need for an explicit mapping between original and synthetic nodes. Consequently, a GM-based method named MCond is introduced, which explicitly learns a sparse mapping matrix to smoothly integrate new nodes into the synthetic graph for inductive representation learning. MCond employs an alternating optimization scheme compared to GCond, allowing the synthetic graph and mapping matrix to take turns updating toward dedicated objectives. Furthermore, CTRL (Zhang et al., 2024a) argues the limited approach of using cosine similarity for gradient matching, leading to biases, and suggests adding gradient magnitudes into the objective function introduced in GCond for a more accurate match. Their empirical findings also show that this approach better aligns frequency distributions between condensed and original graphs.
@@
Despite the effectiveness of the previously mentioned graph condensation methods, Feng et al. (2023) recognize that these methods tend to exhibit fairness issues. By identifying the group fairness [^3] Group fairness in algorithms ensures unbiased and fair treatment across diverse demographic groups. It seeks to prevent any form of discrimination or bias against specific groups within the algorithmic decision-making process (Mehrabi et al., 2021), it demonstrated that as distillation performance increases, fairness (Demographic Parity $\Delta_{DP}$) decreases (Feng et al., 2023). Particularly, it is showcased that, by measuring the fairness of GNNs trained on original graphs versus those trained on condensed graphs, an improvement in performance correlates with heightened fairness issues in the synthetic condensed graph. To address this challenge, FGD is introduced, as a fair graph condensation method. This is achieved by incorporating the coherence metric into the GM loss function outlined in Eq. (8a). Particularly, the coherence metric is a bias calculator that captures the variance of the estimated sensitive group membership. Similarly, to address the fairness issue of current graph condensation methods, Mao et al. (2023) propose graph condensation with Adversarial Regularization (GCARe), which is a method that directly regularizes the condensation process to distill the knowledge of different subgroups fairly into resulting graphs. Also, other studies, such as FedGKD (Pan et al., 2023), utilize gradient matching-based graph condensation in federated graph learning.
@@
where ${\mathcal{C}}$ is the set of node classes, ${\boldsymbol{E}}_{c}$ and $\tilde{\boldsymbol{E}}_{c}$ are the embeddings of nodes with class $c$ in the original and condensed graph, respectively, and $r_{c}$ is the class ratio for class $c$. Likewise, GCDM (Liu et al., 2022) regards the original graph as a distribution of receptive fields and seeks to synthesize a smaller graph whose receptive fields share a similar distribution to that of the original graph. Other works, such as PUMA (Liu et al., 2023d), employ a similar approach for various applications like continual learning.
@@ 3.3.2 Kernel Ridge Regression Methods
where $\epsilon>0$ is a KRR hyper-parameter, $\mathbf{K}_{\mathcal{TS}}$ is the kernel matrix between original and synthetic graphs while $\mathbf{K}_{\mathcal{SS}}$ is the kernel matrix between synthetic graphs [^4] Each kernel indicates infinitely wide multi-layer GNNs trained by gradient descent through squared loss (Du et al., 2019). Also, $\mathbf{y}_{\mathcal{S}}$ and $\mathbf{y}_{\mathcal{T}}$ are the concatenated graph labels from real dataset and synthetic dataset, respectively.
@@ 3.3.3 Other Methods
Also, Zheng et al. (2023b) propose SFGC, a structure-free graph condensation method using a matching-based approach that only outputs the condensed node features $\mathbf{X}^{\prime}$, as the structure information of the real graphs is embedded in ${\bf X}^{\prime}$. Concretely, unlike gradient matching-based methods, SFGC aligns their long-term GNN training trajectories using an offline expert parameter distribution as guidance. Beyond the above efforts, GEOM (Zhang et al., 2024b) makes the first attempt toward lossless graph condensation, i.e., significantly reducing the graph size without any loss in performance. It points out that SFGC (Zheng et al., 2023b) provides biased and limited supervisory signals to difficulty nodes while overlooking easy nodes. To address this, GEOM introduces an expanding window technique to adjust the matching range for difficult and easy nodes, during the process of matching the training trajectories between $\mathcal{T}$ and $\mathcal{S}$. Remarkably, it achieves lossless graph condensation across standard benchmark datasets.
@@ 4.2 Continual Graph Learning
Continual learning  (De Lange et al., 2021) aims to learn on a stream of data from a stationary data distribution, which requires tackling the issue of catastrophic forgetting, i.e., new data can interfere with current knowledge and erase it. The common strategy is memory replay (Kemker and Kanan, 2018), which involves storing representative samples from previous tasks in a buffer to recall knowledge when needed. In the context of graphs, continual learning can be benefited by informative reduced graphs. As introduced in Section 3.3 CaT (Liu et al., 2023c) is applied to continual graph learning by condensing the incoming graph and updating the model with condensed graphs, not the whole incoming graph. To further improve CaT,  Liu et al. (2023d) introduce PUMA which utilizes pseudo-labeling to incorporate data from unlabeled nodes, boosting the informativeness of the memory bank and addressing the problem of neglected unlabeled nodes. In addition, the sparsification method (Zhang et al., 2023b) reduces the number of nodes and edges according to the Ricci curvature evaluation and stores them into a replay buffer for continual learning.
@@ 4.3 Visualization & Explanation
The reduced dataset is not only more accessible for algorithms and computers to parse but also more friendly for people to understand, which leads to the application of visualization and explanation. For instance, Zhao et al. (2018) combine spectral graph coarsening method (Loukas and Vandergheynst, 2018) and sparsification methods (Feng, 2016) to develop a nearly linear time algorithm for multilevel graph visualization. Pyramid Transform (Shuman et al., 2015) selects nodes corresponding to top-$k$ eigenvector repeatedly and creates a multi-resolution view of large graphs. In addition, k-core decomposition has been widely used for graph visualization and finding specific structural characteristics of networks (Alvarez-Hamelin et al., 2005). As mentioned in Section 3.2, some coarsening methods like k-snap (Tian et al., 2008) and CANCEL (Zhang et al., 2010) are tailed for customized discovery, which means the user can set the granularity of the graph structure and then find perspectives they are interested in. Graph condensation is also used for visualization and explanation. For instance, GCond (Jin et al., 2022b) observes patterns from original data by visualizing the reduced graph; GDM (Nian et al., 2023) employs condensation to explain GNN behavior during the training process.
@@ 4.4 Privacy
It has been empirically investigated that reduced datasets offer good privacy preservation (Dong et al., 2022). For example, Bonchi et al. (2014) show random sparsification can achieve meaningful levels of anonymity while preserving features of the original graph.  Arora and Upadhyay (2019) consider edge privacy and provide a mechanism for sparsification of a graph that ensures differential privacy and approximates the spectrum of the original graph. However, the trade-off between the degree of reduction and the utility, i.e., preservation of key information, always exists. Dataset condensation can be a promising technique to solve the dilemma (Dong et al., 2022). In a federated learning framework, where client devices collaboratively contribute to model development by aggregating their updates on a central server, there exists a risk that a malicious server could infer sensitive local information from the model updates sent by these clients. FedGKD (Pan et al., 2023) trains models on condensed local graphs within each client to mitigate the potential leakage of membership of training set (Shokri et al., 2017).
@@ 4.5 Data Augmentation
Graph data augmentation (Zhao et al., 2022, 2021; Zhou et al., 2022) is commonly used to enrich the data and improve the model performance. Methods for graph reduction can be employed to generate various perspectives of a graph by repeatedly applying reductions at different ratios, thereby augmenting the data for subsequent models. For example, HARP (Chen et al., 2018) coarsens a graph in a series of levels and then embeds the hierarchy of graphs from the coarsest one to the original. In each step, HARP initializes node embeddings in the finer-level graph using the mapped embeddings from super nodes in the coarser level. By employing various graphs in each iteration, this method augments the training data. Meanwhile, MILE (Liang et al., 2021) enhances this process by substituting the random walk employed in HARP with GNNs to improve the embedding quality and efficiency. DistMILE (He et al., 2021) advances the MILE framework by adopting high-performance computing techniques. As a condensation method, MSGC (Gao and Wu, 2023) initializes multiple small graphs by various connection schemes and employs gradient matching to optimize them. This process results in a variety of node embeddings sets, increasing diversity and thereby augmenting the data.
@@ 5.4 Distribution Shift
It is a common observation that GNNs often exhibit poor generalization on a test set when there exists a disparity between the distributions of the training and test sets (Wu et al., 2022a; Li et al., 2022). When GNNs are trained on reduced graphs and evaluated on graphs from the original distribution, a distribution shift may occur due to the reduction process, which eliminates substantial amounts of graph elements. However, consensus is lacking regarding the definition of graph data distribution or the selection of specific properties to represent the distribution accurately. While several condensation methods employ a specific type of distribution matching as we mentioned in Section 2, other measures of distribution may change after the reduction, e.g., size shift (Buffelli et al., 2022). Future graph reduction should consider the potential distribution shift issues and preserve distribution-related properties to enhance the generalization of models trained on reduced graphs.
@@ 5.5 Robustness
5.6 Diverse Types of Graphs
Existing graph reduction techniques have primarily concentrated on simple non-attributed or attributed graphs. Nonetheless, graph data in real-world scenarios is increasingly intricate. This complexity is evidenced by the presence of heterophilous graphs (Zheng et al., 2022), heterogeneous graphs  (Bing et al., 2023), directed graphs (Aho et al., 1972), knowledge graphs (Schlichtkrull et al., 2018), hypergraphs (Antelmi et al., 2023) and dynamic graphs (Skarding et al., 2021). Each graph type introduces distinct characteristics, demanding a thorough comprehension of their structures and the adoption of varied graph reduction methods. While some works have developed reduction methods for heterogeneous graphs (Generale et al., 2022), directed graphs (Sugiyama and Sato, 2023) and 3D mesh graphs (Bånkestad et al., 2024), the other types of graphs are yet to be fully explored.